An arrow is shot from a bow at ground level at a speed v0 and an angle of 37o with respect to horizontal. It lands 960 m away (assuming level ground throughout).

Find v0 and max height of arrow
Answer #1

a)

Using expression for max range

R = Vo ^2 sin (2x) / g

960 = Vo^2 sin (2*37) / 9.8

Vo = 98.468 m/s

=========

b)

max hieght is given by

H = ( Vo sin x) ^2/ (2 g)

H = ( 98.468* sin 37)^2/ (19.6)

H = 179.17 m

========)

do comment in case any doubt, will reply for sure.. Goodluck